KjSaif:

Rich deposited money into a bank account that earned 2.5% simple interest each year. After 2 years, he had earned $14.65 in interest on the account. If no other money was deposited into or withdrawn from the account, how much was his initial deposit? Round your answer to the nearest cent.

KjSaif:

@Trillboij

KjSaif:

oh your here ._.

Trillboij:

lol

KjSaif:

what do you guys think? @Trillboij @wo1f0mon @Ana23 @Awolflover1 @ssddss02 @studying

KjSaif:

@Reynelle

Trillboij:

I'm not sure.

KjSaif:

:(

KjSaif:

ok

wo1f0mon:

answer choices?

KjSaif:

It's infinite

KjSaif:

No choices :(

KjSaif:

@wo1f0mon

wo1f0mon:

i cant help sowe

KjSaif:

it's cool @wo1f0mon

KjSaif:

:(

KjSaif:

i'm thinking $0.37

KjSaif:

im so lost though......

KjSaif:

@Korianna @Bob

Reynelle:

I can help

Trillboij:

O:

KjSaif:

maybe it's $0.40

Reynelle:

trill what?

Trillboij:

I was thinking $0.48

Trillboij:

Nothing rey. :3

Reynelle:

k den

Trillboij:

;)

Reynelle:

I will try to answer

KjSaif:

ok @Reynelle

KjSaif:

just googled and found this....

KjSaif:

Calculate Total Amount Accrued (Principal + Interest) A = P(1 + rt) Calculate Principal Amount, solve for P P = A / (1 + rt) Calculate rate of interest in decimal, solve for r r = (1/t)(A/P - 1) Calculate rate of interest in percent R = r * 100 Calculate time, solve for t t = (1/r)(A/P - 1)

Reynelle:

Does that help you?

Trillboij:

That's the formula

Trillboij:

right?

Reynelle:

i gtg i will try to get answer n the morning :3

KjSaif:

I GOT IT!!!!

Reynelle:

GOOD!!!!!

Reynelle:

alright byes

KjSaif:

$293

KjSaif:

WAIT

KjSaif:

thats for the first year.....................

KjSaif:

-_-

KjSaif:

wait no it's not because the rate was 14.65

KjSaif:

2.5%

KjSaif:

find a random number that equals the first years interest rate

KjSaif:

(7.325)

KjSaif:

so 293 divided by 2.5 = 7.325

KjSaif:

Can you guys review my 5 questions?

CloroxBleach:

e.e

umer:

e-e

MsBrains:

^&^